OpenStudy (anonymous):

Identify the excluded value for the rational function. y=4x-1 PLEASE help me understand this!!!

OpenStudy (solomonzelman):

this function here, that you posted is a polynomial

OpenStudy (solomonzelman):

that function includes any x-value in it's domain. It is continuous over \((-\infty,~+\infty)\)

OpenStudy (solomonzelman):

there are no excluded x values in a polynomial.

OpenStudy (anonymous):

Are you sure because I have already tried answering this and I put 0, and it says that i'm wrong. The options are: 0, 1, 2, and 3.

OpenStudy (solomonzelman):

you put y=4x-1, you see that, correct?

OpenStudy (solomonzelman):

maybe it is \(\large\color{black}{ \displaystyle y=\frac{\rm something}{\rm another~~something } }\)

OpenStudy (anonymous):

ya it's y=4/x-1

OpenStudy (solomonzelman):

oh, \(\large\color{black}{ \displaystyle y=\frac{4}{x-1} }\)

OpenStudy (solomonzelman):

that is much better:)

OpenStudy (solomonzelman):

what number can't you divide by, do you know?

OpenStudy (anonymous):

x?

OpenStudy (solomonzelman):

4 ÷ what = undefined or indeterminate ?

OpenStudy (solomonzelman):

you can divide by 1, you can divide by 3, you can divide by -0.0003, but you can NOT divide by 0. right?

OpenStudy (solomonzelman):

you can divide by anything, but not by 0

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

so, in your function \(\large\color{black}{ \displaystyle y=\frac{4}{x-1} }\) , there is one x-value that will not give you an adequate output, because you would need to divide by 0.

OpenStudy (solomonzelman):

you are dividing by x-1. so, when (for what x-value) is this the same as dividing by 0 ?

OpenStudy (solomonzelman):

(in other words, when "x-1" is equivalent to 0? when x=what ?)

OpenStudy (anonymous):

undefined?

OpenStudy (solomonzelman):

ok, lets go about it in a different way...

OpenStudy (solomonzelman):

you can't divde by 0, you know that?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

in our case we are dividing 4 by "x-1", yes?

OpenStudy (solomonzelman):

now, when dividing by x-1, is going to be the same thing as dividing by 0, that will give us an undefined result.

OpenStudy (solomonzelman):

dividing by x-1 is same as dividing by 0, when x-1 is same as 0. there is one x-value for which x-1 is same as 0. x-1=0 when is x-1=0?

OpenStudy (anonymous):

idk

OpenStudy (solomonzelman):

I tried my best, I apologize, but I don't think I can help you more. I am capable of you the answer, but I am not going to.

OpenStudy (solomonzelman):

of giving you*

OpenStudy (solomonzelman):

bye

OpenStudy (anonymous):

Well thanks anyway

OpenStudy (solomonzelman):

-;( yw

OpenStudy (anonymous):

@SolomonZelman I think I am understanding this more. x-1=0 is 1-1=0 is this correct?

OpenStudy (anonymous):

So this would mean that the excluded value is 1 right? @SolomonZelman

OpenStudy (solomonzelman):

yes, yes very good!

OpenStudy (solomonzelman):

the excluded value is 1. x≠1

OpenStudy (solomonzelman):

:)

OpenStudy (anonymous):

Thank you so much for helping!!